The difference between two numbers is 18. The sum of twice the smaller number and three times the larger number is 784. What are the numbers?

Answer by checkley79(3341)

X-Y=18
X=18+Y
2Y+3X=784
2Y+3(18+Y)=784
2Y+54+3Y=784
5Y=784-54
5Y=730
Y=730/5
Y=146 ANS FOR THE SMALLER NUMBER.
X-146=18
X=18+146
X=164 ANS. FOR THE LARGER NUMBER.
PROOF:
2*146=3*164=784
292+492=784
784=784